OpenStudy (anonymous):

Choose the equivalent system of linear equations that will produce the same solution as the one given below. 6x + 2y = -6 3x - 4y = -18 A. 12x + 4y = -12 15x = -30 B. 8x + 4y = -4 14x = -10 C. 6x - 8y = -36 -6y = -42 D. 6x - y = -15 3y = 9

jimthompson5910 (jim_thompson5910):

What is the solution to this system? 6x + 2y = -6 3x - 4y = -18

OpenStudy (anonymous):

Idk how to solve them..

jimthompson5910 (jim_thompson5910):

Check out this page http://coolmath.com/algebra/12-2x2-systems-of-equations/03-solving-by-elimination-addition-01.htm

OpenStudy (anonymous):

Okay, so do I do all of that for them individually or as if the bottom was the second line?

jimthompson5910 (jim_thompson5910):

you solve each system and see which solutions match with the solutions of 6x + 2y = -6 3x - 4y = -18

OpenStudy (anonymous):

Oh, lord. Okay.